OpenStudy (anonymous):

Tickets to a concert cost $2 for children and $5 for adults. If 630 people attended the concert and the ticket sales totaled $2,910, how many children attended the concert? 36 50 80 120

OpenStudy (anonymous):

Do u still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which one do U think it is?

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

a = amount of adults c = amount of children 5 bucks per adult, sales for that will be 5 * a, or 5a 2 bucks per child, sales for that will be 2 * c, or 2c so 630 people attended, adults and children summed up will be a+c=630 so total sales was $2,910 so adult sales and children sales summed up will be 5a+2c=2910 ------------------------------------------------ from a + c = 630 we could solve for "a" and say that a=630−c so now we know that a = 630-c, so let us use that in the 2nd equation 5(630−c)+2c=2910 solve for "c" to find out how many children atended

OpenStudy (anonymous):

@ladybug99 please don't give direct answers this is a violation of the openstudy code of conduct. This is a warning. :)

OpenStudy (anonymous):

^^ and who are you ??

OpenStudy (anonymous):

What do u mean?

OpenStudy (anonymous):

what do you mean violation to openstudy code like if i wanted to tell the answer i can

OpenStudy (anonymous):

We just want them to learn rather than just get the answer. They won't learn anything.

OpenStudy (anonymous):

If u would like I could refer u to the code of conduct for openstudy.

OpenStudy (anonymous):

i mean they can ask how to get the answer but if you dont like it report it okaii

OpenStudy (anonymous):

We just don't want to give direct answer u have to let the asker have some involvement. I will not report u I dot think u can do that. Just please don't give direct answers.

OpenStudy (anonymous):

U simply blocked me from following u. Thank u and have a nice day! :)